OpenStudy (anonymous):

Identify the number that does not belong with the other three. Explain your reasoning. 50.1 repeating 1, negative 50 over 2, negative 50.1, square root 50

OpenStudy (michele_laino):

hint: \[50.\bar 1 = \frac{{501 - 50}}{9} = \frac{{451}}{9}\]

OpenStudy (michele_laino):

\[\frac{{ - 50}}{2} = - \frac{{25}}{1}\]

OpenStudy (michele_laino):

\[ - 50.1 = - \frac{{501}}{{10}}\]

OpenStudy (anonymous):

what do you mean? @Michele_Laino

OpenStudy (michele_laino):

\[\sqrt {50} = 5\sqrt 2 \]

OpenStudy (michele_laino):

and: \[\sqrt 2 \] can not be express as a ratio between two integer numbers

OpenStudy (anonymous):

so is it square root 50

OpenStudy (michele_laino):

yes! It is an irrational number

OpenStudy (anonymous):

thank you sooo much!!

OpenStudy (michele_laino):

thank you!! :)